# Exchangeability in Neural Network Architectures and its Application to Dynamic Pruning

### v1

@@ Abstract
Neural networks (NNs) are equipped with increasingly many parameters and require more and more resource for deployment. Researchers have explored various ways to improve the efficiency of NNs by identifying and reducing the redundancy, such as pruning or quantizing unimportant weights. Symmetry in the NN architectures has been identified by prior work as a possible type of redundancy, but exploiting it for efficient inference is not yet explored. In this work, we formalize the symmetry of parameters and intermediate values in NNs using the statistical property of exchangeablility. We identify that exchangeable values in NN computation may contain overlapping information, leading to redundancy. Exploiting the insight, we derive a principled general dynamic pruning algorithm ExPrune to remove symmetry-induced redundancy on a per-input basis. We also provide an instantiation of ExPrune that performs neuron-level dynamic pruning by predicting negative inputs to ReLU activations. We evaluate ExPrune on two computer vision models, one graph model and one language model. ExPrune provides 10.98–26.3% reduction in FLOPs with negligible accuracy drop and 21.01–39.05% reduction in FLOPs with at most 1% accuracy drop. We also demonstrate that ExPrune composes with static pruning. On models that have been aggressively pruned statically, ExPrune provides additional 10.24–11.11% reduction in FLOPs with negligible accuracy drop and 13.91–14.39% reduction in FLOPs with at most 1% accuracy drop.
@@ 1 Introduction
Deep learning has achieved tremendous success in the past few years, with applications in computer vision visiontransformers ; ResNet , natural language processing word2vec ; EdgeBERT , graph processing WeihuaGCN ; GCNWelling ; EnEquivariantGNN , generative AI gpt2 ; llama ; OPT . The demand for increasingly large and powerful deep neural networks (NNs) greatly increased the memory/compute footprint of NN inference tasks DynamicNNSurvey ; modelcompressionsurvey . To tame resource usage, researchers have developed a number of NN optimizations that statically reduce model size before deployment, including static pruning SongPruning ; SurveyPruning , quantization saha2024compressing ; QuantizationSurvey ; YoshuaBNN ; BinaryNNSurvey , knowledge distillation HintonDistillation ; SurveyDistillation , and neural architecture search NASsurvey . To a lesser extent, researchers have also developed optimizations that dynamically prune the inference computation at runtime on a per-input basis MixtureOfExpert ; BranchyNet ; snapea . A common thread linking many of these methods is that they focus on eliminating redundancies present in the trained model.
@@
In this work, we present ExPrune, a general, dynamic pruning optimization that enables multi-granularity (e.g., neuron/kernel/layer) partial computation on a per-input basis. This optimization capitalizes on the presence of exchangeable model parameters and intermediate values, a property that we show is implied by certain forms of symmetry. Exchangeability is a statistical property that implies identically distributed and symmetric interdependence among random variables dean1990linear . Intuitively, exchangeable model parameters and values contain overlapping information, which induces redundant computation. In this work, we show that certain groups of parameters are trained to be exchangeable with respect to random initializations and thus produce exchangeable intermediate values – these quantities are exchangeable by construction. We also establish a formal link between exchangeability and permutation invariance of parameters, a related model property that has not yet been exploited for efficiency. Because the ExPrune is grounded in these theoretical results, it can generalize across model architectures, and we identify exchangeable parameter/value patterns in a range of modern NNs, such as Convolution Neural Networks (CNNs), Graph Neural Networks (GNNs) and transformer-based language models (LMs).
To our knowledge, this work is both the first to model NN symmetry with exchangeability and to use symmetry-induced redundancies to perform dynamic pruning. Theoreticians have previously linked NN symmetries to redundancy EmpiricalImpactSymmetry and studied the impact of these symmetries on NN models’ ability to generalize SharpMinimaGeneralize , their interpretability SymmetryInternalRepresentation ; symmetryDNNinternalrepresentations , and loss landscape SymmetryFlatMinimaConservedQuantities . These works did not link the symmetry to exchangeability, or mechanized it to perform static/dynamic pruning.
@@
To demonstrate the efficacy of this method, we instantiate the ExPrune algorithm to prune computations of each neuron’s activation by predicting negative inputs to the ReLU activation function. We evaluate this ExPrune instantiation on CNNs, GNNs, and LMs. We find that ExPrune can reduce the inference cost of these NNs by 10.98–26.3% with negligible drop on model accuracy. We also empirically validate that ExPrune can compose with static pruning, providing additional efficiency improvements.
Contributions: We present the following contributions:
@@
  • Formalization of NN symmetries with statistical exchangeability (Section 3): To the best of our knowledge, this is the first work to formalize the symmetry in NN architectures using statistical exchangeability. The formalism provides insights into symmetry-induced redundancy.
@@
  • Dynamic pruning algorithm for reducing symmetry-induced redundancy (Section 4): We present ExPrune, a general dynamic pruning algorithm that removes symmetry-induced redundancy. The algorithm is based on our theoretical formulation, can be applied to various NN architectures, and composes with most existing approaches for efficient inference. We also present an instantiation of ExPrune that does neuron-level dynamic pruning by performing negative prediction of inputs to ReLU.
@@
  • Evaluation (Section 5): We evaluate ExPrune on CNNs, GNNs, and LMs. We demonstrate that ExPrune provides 10.98–26.3% reduction in FLOPs with negligible accuracy drop and 21.01–39.05% reduction in FLOPs with at most 1% accuracy drop. We also empirically demonstrate that ExPrune composes with static pruning. On models that have been aggressively pruned statically, ExPrune provides additional 10.24–11.11% reduction in FLOPs with negligible accuracy drop and 13.91–14.39% reduction in FLOPs with at most 1% accuracy drop. Our evaluation code is released at https://anonymous.4open.science/r/Exchangeable-NN-FBC7.
2 Related Work
Symmetry in Deep Learning. Symmetry in NNs and its impact have been extensively studied by theoreticians. Symmetry is known to affect model generalization SharpMinimaGeneralize , the loss landscape SymmetryFlatMinimaConservedQuantities ; EmpiricalImpactSymmetry , interpretability SymmetryInternalRepresentation ; symmetryDNNinternalrepresentations , Bayesian NN inference DetrimentalInvarianceBayesianNN , graph metanetwork processing GraphMetanetwork , and aligning symmetric sub-structures helps improving efficiency and performance of ensemble NN models  ModelFusion ; transformerfusion ; GitReBasin . Researchers have also explored deliberately breaking the symmetry in NN architectures EmpiricalImpactSymmetry ; improvingidentifiabilityNNbayesianinference ; removingsymmetriesrevealstructureofflatregionsoflandscape ; SymmetryAwareBNN ; RemoveSymmetriesControlExpressivity . A related but different concept is equivariance DeepSets ; EnEquivariantGNN ; GroupEquivariantCNN , a feature of special NN architectures that the NNs produce consistent outputs under symmetry transformations of the inputs. To the best of our knowledge, we are the first to use dynamic pruning to reduce symmetry-induced redundancy.
Exchangeability in Deep Learning. Statistical exchangeability has various applications in deep learning, but prior work focused on exchangeable data. Observing exchangeability of certain input data such as sets, special NN architectures have been proposed to process them LikelihoodFreeExchangeableNN ; BRUNORucurrentModelForExchangeableData ; ProbabilisticSymmetriesAndInvariantNN ; MCMCsamplingBayesianNN . Conformal prediction provides a prediction set with guaranteed error rate for any NN model assuming exchangeability of data sequence fontana2023conformal ; kuchibhotla2020exchangeability . To the best of our knowledge, this is the first work to link statistical exchangeability to the symmetry in NN architecture.
Static NN Model Optimizations. Various techniques have been proposed to derive efficient NN models with smaller sizes and fewer parameters, including pruning SongPruning ; SurveyPruning , quantization saha2024compressing ; QuantizationSurvey ; YoshuaBNN ; BinaryNNSurvey , knowledge distillation HintonDistillation ; SurveyDistillation , and neural architectural search NASsurvey . These methods are statically applied before model deployment. In contrast, ours is dynamic and applies on a per-input basis during inference. In our evaluation (Section 5.2), we show that our method composes with static pruning.
Dynamic Pruning at Inference. Researchers have explored coarse-grained dynamic pruning methods, which are applied on a per-input basis during NN inference SurveyPruning . They explored dynamically pruning layers BranchyNet ; EdgeBERT ; DynamicNNSurvey , tokens dynamiccontextpruneLLM , channels RuntimeNeuralPruning ; elkerdawy2022fire , spatial domain FrequencyDomainPrune , and branches MixtureOfExpert . These methods are specialized to certain model architectures and coarse-grained, operating on structures larger than neurons. In contrast, our method can operate at very fine granularity (neuron level) and can be generalized across multiple architectures and granularities.
SnaPEA snapea and ComPreEND ComPreEND explored neuron-level dynamic pruning for convolution. They reordered the computation so that the partial sum first monotonically increases then decreases, and terminate computation when the partial sum is negative while monotonically decreasing. These methods are narrow in scope, focusing on CNNs and do not natively support models with normalization layers BatchNorm ; LayerNorm or skip connections ResNet in modern CNNs. These also do not generalize to other architectures, because they assume that the input values are positive, which may not be true, e.g., in transformers attention and GNNs GCNWelling . In contrast, our method supports shortcut connections and normalization layers, and works with many model architectures.
Prior work video_mlsys ; convrelu++ exploited similar patches in input feature maps to derive upper bound of the weighted sum given the partial sum, and terminated early when the upper bound is below zero. These methods only take effect on similar input patches and are specialized to CNNs and video processing. In contrast, our method works with many model architectures, and does not require similar input patches. We note that our method can compose with theirs as the confidence metric in our system ExPrune (Section 4) can incorporate their upper bound information.
@@ 3 Exchangeability in Neural Networks
Figure 1: Illustration of exchangeable parameters and values in ReLU-activated MLP without bias. $a,b,c$ are neuron activations. $W^{\prime}$ and $W$ are weight matrices. Symmetric parameters and values with different colors are exchangeable, thus identically distributed.
We present the background and related theories in Section 3.1, our theories on exchangeability in NNs in Section 3.2, and then identify exchangeable parameters and values in popular NN architectures 3.3. We use a simple ReLU-activated multi-layer-perceptron (MLP) without bias as a running example to provide intuitions, shown in Figure 1. In this example, we focus on two hidden layers (colored) with weights $W^{\prime}$ and $W$. Denote the neuron activations of sequential layers shown as $a,b,c$. We have $b=\mathrm{ReLU}(W^{\prime}a),c=\mathrm{ReLU}(Wb)=\mathrm{ReLU}(W\mathrm{ReLU}(W^{\prime}a))$. Note that the choice of ReLU is immaterial in this section, as our theorems are independent of specific activation functions.
An obvious symmetry in the model is that permuting the hidden (colored) neurons’ positions (the connected edges move along) does not change the function mapping from $a$ to $c$. We delineate the symmetry using a statistical property called exchangeability. There are two forms of exchangeability in Figure 1: exchangeable parameters are the colored edges $\zeta{}_{i}=(W^{\prime}_{1i},W^{\prime}_{2i},W^{\prime}_{3i},W_{i1},W_{i2})$, and exchangeable values are colored intermediate values $\xi{}_{i}=b_{i}$ or $\xi{}_{i}=W_{1i}b_{i}$. Intuitively, exchangeability here means that if we view $\zeta{}_{i}$’s ($i=1,2,3,4$) as random variables, with respect to random initializations of the NNs, they are from the same distribution and have symmetric interdependence relationship. The same is true for $\xi{}_{i}$’s, for any given input to the NN. We next introduce statistical exchangeability formally, and relevant definitions and theories in prior work.
@@ 3.1 Background: Statistical Exchangeability and Parameter Space Symmetry
We first define statistical exchangeability. Let $P$ be a $n\times n$ permutation matrix for some $n$. The following definition is taken from Kuchibhotla kuchibhotla2020exchangeability (assuming the probability density function exists).
Definition 1 (Exchangeablility).
Suppose $\zeta{}=\left(\zeta{}_{1},\ldots,\zeta{}_{n}\right)\in\mathcal{X}^{n}$ is a vector of random variables, $\zeta{}_{i}$’s are exchangeable iff their joint probability density function $p(\zeta{})$ is invariant to input permutations, i.e., $\forall$ permutation matrix $P$ and $\zeta{}\in\mathcal{X}^{n}$, $p(P\zeta{})=p(\zeta{})$.
@@
The following theorem states a condition under which a transformation on exchangeable random variables preserves their exchangeability kuchibhotla2020exchangeability ; dean1990linear (stronger condition).
@@
Suppose $\zeta{}=\left(\zeta{}_{1},\ldots,\zeta{}_{n}\right)\in\mathcal{X}^{n}$ is a vector of exchangeable random variables. Fix a transformation $G:\mathcal{X}^{n}\rightarrow\mathcal{X}^{n}$. If $G$ is permutation equivariant, i.e., $\forall$ permutation matrix $P$ and $\zeta{}\in\mathcal{X}^{n}$, $PG(\zeta{})=G(P\zeta{})$, then $G(\cdot)$ preserves exchangeability of $\zeta{}$.
Parameter space symmetry is a relevant concept to describe the symmetry in parameter space. Given a NN architecture with $N$ real-valued parameters, we denote the NN function parameterized by $\theta{}\in\mathcal{R}^{N}$ as $f_{\theta{}}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{X},\mathcal{Y}$ are input and output spaces respectively. A parameter space symmetry is defined as follows EmpiricalImpactSymmetry .
@@
In the example (Figure 1), each permutation on $\zeta{}_{i}=(W^{\prime}_{1i},W^{\prime}_{2i},W^{\prime}_{3i},W_{i1},W_{i2})$ (the associated weights of each hidden neuron) is a parameter space symmetry.
@@ 3.2 Exchangeable Parameters and Values in Neural Networks
In this section, we prove that (1) for a group of parameters $\zeta{}_{i}$’s, if every permutation of them is a parameter space symmetry, they are exchangeable after training, with respect to random initializations (Theorem 2), and (2) if a group of values $\xi{}_{1},\xi{}_{2},\ldots,\xi{}_{n}$ are computed with $\zeta{}_{1},\zeta{}_{2},\ldots,\zeta{}_{n}$, respectively, they are exchangeable as well (Theorem 3). Below is our key insight in the proof.
[…]
@@
Given an NN architecture with function $f_{\theta{}}$, assume that $\zeta{}_{i}$’s are initialized exchangeable. If for any permutation matrix $P$, $\omega_{P}$ defined above is a parameter space symmetry, then $\zeta{}_{i}$’s are exchangeable after training, with respect to random initializations.
@@
Here we assume using simple gradient descent for supervised learning on classification tasks for simplicity, but the the proof easily generalizes to other optimization algorithms. As we are only interested in $\zeta{}_{i}$’s, we formalize a training step as a transformation on only the $\zeta{}_{i}$’s. [^1] In fact, one can prove that permuting $\zeta{}_{i}$’s does not affect the update to other parameters in $\theta{}^{\prime}$. Denote the NN loss function as $L_{\zeta{}}:\mathcal{X}\times\mathcal{Y}\times\mathcal{R}^{N-mn}\rightarrow\mathcal{R}$, parameterized by $\zeta{}$. $L_{\zeta{}}(x,y,\theta{}^{\prime})=\psi(f_{\theta{}}(x),y)$ for some metric function $\psi$ such as cross entropy. A training step $G_{S}:(\mathcal{R}^{m})^{n}\rightarrow(\mathcal{R}^{m})^{n}$ takes a training batch of $B$ samples $S\in(\mathcal{X}\times\mathcal{Y})^{B}$, and does a gradient descent step $G_{S}(\zeta{})=\zeta{}-\gamma\sum_{(x,y)\in S}\nabla_{\zeta{}}L_{\zeta{}}(x,y,\theta{}^{\prime})$, where $\gamma$ is the learning rate, and $\nabla_{\zeta{}}L_{\zeta{}}(x,y,\theta{}^{\prime})$ is the gradient of $L_{\zeta{}}$ with respective to $\zeta{}$, evaluated at $(x,y,\theta{}^{\prime})$.
@@
Note that $\nabla_{P\zeta{}}L_{P\zeta{}}$ is not the same as $\nabla_{\zeta{}}L_{P\zeta{}}$, as the derivatives in the gradient vector of the former case shall match the parameter order in $P\zeta{}$ rather than $\zeta{}$. Since $\omega_{P}$ is a parameter space symmetry, by definition 2, $\forall\zeta{}\in(\mathcal{R}^{m})^{n}$, $f_{\theta{}}$ and $f_{\omega_{P}(\theta{})}$ are the same function, and thus $L_{\zeta{}}$ and $L_{P\zeta{}}$ are the same function. Therefore, it is straightforward that $\nabla_{P\zeta{}}L_{P\zeta{}}(x,y,\theta{}^{\prime})=P\nabla_{\zeta{}}L_{\zeta{}}(x,y,\theta{}^{\prime})$ for all $x,y,\zeta{},\theta{}^{\prime}$. This implies that $PG_{S}(\zeta{})=G_{S}(P\zeta{})$. ∎
@@
Under the conditions of Theorem 2, after training, for any NN input $x\in\mathcal{X}$, define $\xi{}_{i}$’s as $\xi{}_{i}=g_{\theta{}^{\prime},\zeta{}_{i}}(x)$, for some function $g$ that can be parameterized by $\theta{}^{\prime}$ and one $\zeta{}_{i}$. $\xi{}_{i}$’s are exchangeable with respect to random NN initializations.
@@
Since $\theta{}^{\prime}$ is shared among $\xi{}_{i}$’s, the only variable in $\xi{}_{i}$ is $\zeta{}_{i}$. $(\zeta{}_{1},\ldots,\zeta{}_{n})\rightarrow(\xi{}_{1},\ldots,\xi{}_{n})$ is an exchangeability preserving transformation by Theorem 1. Therefore, $\xi{}_{i}$’s are exchangeable. ∎
Note that it is possible that $g_{\theta{}^{\prime},\zeta{}_{i}}$ uses all or only a part of $\zeta{}_{i}$ in its computation. Given a group of exchangeable parameters $\zeta{}_{i}$’s, there might be multiple groups of exchangeable values $\xi{}_{i}$’s in NN computation, each with a different $g$ function.

### v2

# Exchangeability in Neural Networks and its Application to Dynamic Pruning

@@ Abstract
Modern neural networks (NN) contain an ever-growing number of parameters, substantially increasing the memory and computational cost of inference. Researchers have explored various ways to reduce the inference cost of NNs by reducing the model size before deployment and dynamically pruning the inference computation at runtime. In this work, we present ExPrune, a general, dynamic pruning optimization that enables multi-granularity partial computation on a per-input basis. ExPrune requires no change to the model architecture or the training algorithm. ExPrune is based on our theoretical results that the relationship between certain model parameters and intermediate values can be described by a statistical property called exchangeability. By identifying exchangeable parameters and values in the model, we are able to first partially evaluate the network, analyze the statistics of the partial results, and make pruning decisions on the fly. Because ExPrune is theory grounded, it generalizes across model architectures in different problem domains. We evaluate ExPrune on one computer vision models, one graph model and one language model. ExPrune provides 10.98–17.33% reduction in FLOPs with negligible accuracy drop and 21.61–27.16% reduction in FLOPs with at most 1% accuracy drop. We also demonstrate that ExPrune composes with static magnitude pruning. On models that have been aggressively statically pruned, ExPrune still provides additional 10.24–11.11% reduction in FLOPs with negligible accuracy drop and 13.91–14.39% reduction in FLOPs with at most 1% accuracy drop.
@@ 1 Introduction
Modern neural networks (NN) contain an ever-growing number of parameters, substantially increasing the memory and computational cost of inference (DynamicNNSurvey, ). To tame resource usage, researchers have developed a number of NN optimizations that statically reduce model size before deployment, including static pruning (SurveyPruning, ), quantization (saha2024compressing, ), knowledge distillation (SurveyDistillation, ), and neural architecture search (NASsurvey, ). To a lesser extent, researchers have also developed optimizations that dynamically prune the inference computation (BranchyNet, ; elkerdawy2022fire, ). Dynamic pruning is a runtime optimization that allows parts of the inference computation to be skipped on a per-input basis, enabling partial computation and improving performance. Prior dynamic pruning methods are usually specialized to specific model architectures and network granularities, and often require changes of the model architectures and the training algorithms. These limitations make them harder to apply to a broad class of models and training strategies.
@@
In this work, we present ExPrune, a general, dynamic pruning optimization that enables multi-granularity (e.g., neuron/kernel/layer) partial computation on a per-input basis. ExPrune requires no changes to the model architecture or the training algorithm, as it exploits structures already present in the model. Specifically, ExPrune capitalizes on the presence of exchangeable model parameters and intermediate values. Exchangeability is a statistical property that implies identical distribution and symmetric interdependence among random variables (dean1990linear, ). By formalizing model training as drawing a random model from a distribution of models with respect to random initialization, we prove that certain model parameters have exchangeable marginal distributions, and so do the intermediate values computed with them. The property of exchangeability enables us to first partially evaluate the network, analyze the statistics of the partial results, and prune some computation on the fly. Because the ExPrune is grounded in these theoretical results, it can generalize across model architectures. We identify exchangeable parameter/value patterns in a range of modern NNs, including Convolution Neural Networks (CNNs), Graph Neural Networks (GNNs) and transformer-based language models (LMs).
@@
To demonstrate the efficacy of this method, we instantiate the ExPrune algorithm to prune two common structures present in NNs. We evaluate ExPrune on CNNs, GNNs, and LMs. We find that ExPrune can reduce the inference cost of these NNs by 10.98–17.33% with negligible drop on model accuracy. ExPrune is more general and outperforms the most similar prior work while requiring much less branching operations. We also empirically validate that ExPrune can compose with static pruning, providing additional efficiency improvements.
Roadmap and Contributions: Section 2 presents a simple running example, providing intuitions of our theory and algorithm. Section 3 and Section 4 present the formalism and details. Section 5 presents our evaluation. We present the following contributions:
@@
  • Analysis of statistical exchangeability in NNs (Section 3): To our knowledge, this work is both the first to model the relationship among NN parameters with exchangeability and to perform dynamic pruning exploiting this property. The formalism provides insights into symmetry-induced redundancy identified by theoreticians (EmpiricalImpactSymmetry, ).
@@
  • Dynamic pruning algorithm (Section 4): We present ExPrune, a general dynamic pruning algorithm based on our theoretical formulation. ExPrune can be applied to various NN architectures at various network granularities, and can compose with existing approaches for efficient inference. We also present two instantiations of ExPrune that prunes two common structures in NNs.
@@
  • Evaluation (Section 5): We evaluate ExPrune on CNNs, GNNs, and LMs. We demonstrate that ExPrune provides 10.98–17.33% reduction in FLOPs with negligible accuracy drop and 21.61–27.16% reduction in FLOPs with <1% accuracy drop. We also demonstrate that ExPrune composes with static magnitude pruning. On models that have been aggressively statically pruned, ExPrune provides additional 10.24–11.11% reduction in FLOPs with negligible accuracy drop and 13.91–14.39% reduction in FLOPs with at most 1% accuracy drop.
2 Running Example: Dynamic Pruning of ReLU-activated MLP
Figure 1: Dynamic pruning with ExPrune algorithm, grounded by our theory of exchangeability.
Figure 2: MLP without bias. $a,b,c$ are neuron activations. $W^{\prime}$ and $W$ are weight matrices. Parameters and values with the same color are exchangeable, thus identically distributed.
Figure 3: Distributions of weights $W^{\prime}_{i,1}$’s, $W_{2,i}$’s, and values $b_{i}$’s on one input over 500 trained models.
In this section, we show how we use ExPrune to dynamically prune (Figure 1) a simple ReLU-activated multi-layer-perceptron (MLP) without bias (Figure 3), and the theoretical groundings behind it. In this example, we focus on two hidden layers (colored) with weights $W^{\prime}$ and $W$. Denote the neuron activations of sequential layers shown as $a,b,c$. We have $b=\mathrm{ReLU}(W^{\prime}a),c=\mathrm{ReLU}(Wb)=\mathrm{ReLU}(W~\mathrm{ReLU}(W^{\prime}a))$.
Dynamic pruning. Figure 1 shows an illustration of our ExPrune dynamic pruning algorithm. It prunes the computation at runtime on a per-input basis. Assume that we are currently computing $c_{1}=\mathrm{ReLU}(\sum_{i=0}^{n}W_{1,i}b_{i})$ at inference. Instead of computing all $n$ terms $W_{1,i}b_{i}$ for $i\leq n$, ExPrune first computes $k$ terms, and decides whether computation of other terms can be skipped based on these $k$ terms. In this example, if $\sum_{i=0}^{k}W_{1,i}b_{i}\ll 0$, we can predict that the sum of all $N$ terms is likely to be negative, so $c_{1}$ likely equals 0, and we can prune the computation of the other terms. We can predict the negativity of the final sum by simply comparing the partial sum with a threshold, or using more sophisticated methods like conducting a statistical test on the computed terms. This dynamic pruning techniques saves multiplication operations, translating to reduced compute and data movement.
In the above optimization, we use a partial sum to estimate the result of the full sum. We therefore implicitly assume that the terms $W_{1,i}b_{i}$’s can be regarded as samples drawn from the same distribution. Only if this assumption holds can we infer information about the other terms from the computed $k$ terms, enabling us to approximate the full result with partial computation. We develop a theory, which proves that these terms have exchangeable distributions, a stronger condition than the above assumption, because it in addition entails that their statistical dependence among each other is symmetric.
Training as drawing a random model. Our theory is grounded in a statistical view of training as a random process, where the statistical properties of parameters are derived over the distribution of trained models. At the beginning of training, parameters are randomly initialized – this serves as the source of randomness in the training algorithm. Each random instantiation of the initialized model yields a different trained model. By analyzing all possibilities, the trained model can be viewed as a sample from a distribution of models, with respect to the random initializations. The distribution of the trained models, and the marginal distribution of any parameter in the trained models, are highly complex and intractable to derive precisely. Nonetheless, we can derive that certain parameters have exchangeable marginal distributions, and so do values computed from them.
Exchangeable parameters. We develop a theory that analyzes the structure of the model and find that certain weights have exchangeable marginal distributions. For the MLP example, we analytically derive that for a fixed $j$ and different $i$’s, $W^{\prime}_{i,j}$’s have exchangeable distributions, and so do $W_{j,i}$’s (as shown in Figure 3). To illustrate, we empirically study the identical distribution property, a weaker property entailed by exchangeability in Figure 3. We train 500 NNs (2-layer MLPs with 128 hidden neurons that classify MNIST) from 500 random initializations, and plot over the 500 trained models, the distribution of each $W^{\prime}_{i,1}$, and the distribution of each $W_{2,i}$, for 128 different $i$’s in Figure 3. We can clearly see that $W^{\prime}_{i,1}$’s are identically distributed, and $W_{2,i}$’s also follow an identical distribution, which is different from the distribution of $W^{\prime}_{i,1}$’s.
Exchangeable values. If we know certain parameters have exchangeable distributions, the values computed from these parameters will also have exchangeable distributions, given any input. In the example (Figure 3), $b_{i}$’s have exchangeable distributions, and so do the terms $W_{1,i}b_{i}$’s. In Figure 3, we also plot the distribution of each $b_{i}$ normalized to [1,1][-1,1] on a specific MNIST image over 500 trained models to show that $b_{i}$’s are identically distributed. A large density of the distribution is near the minimum because $b_{i}$’s are ReLU activated and often 0. Specifically, the theoretical result that different terms $W_{1,i}b_{i}$’s can be viewed as samples from the same distribution grounds the dynamic pruning algorithm ExPrune.
@@ 3 Exchangeability in Neural Networks
We formally define and derive exchangeability properties in the trained model. Note that this section is a pure theoretical analysis and does not make any change to the training algorithm or NN architectures.
@@ 3.1 Background: Statistical Exchangeability and Parameter Space Symmetry
Let $P$ be a $n\times n$ permutation matrix for some positive integer $n$. The following definition of exchangeability is taken from prior work kuchibhotla2020exchangeability (assuming the probability density function exists).
Definition 1 (Exchangeability).
@@
Exchangeability is a stronger condition than identical distribution, as it in addition implies the symmetric interdependence among $\zeta{}_{i}$’s, but it is weaker than $\mathit{iid}$ (probabilitytheorybook, ). The following theorem states a condition under which a transformation on exchangeable random variables preserves their exchangeability (kuchibhotla2020exchangeability, ; dean1990linear, ) (stronger condition).
@@
[…]
Given a NN architecture with $N$ real-valued parameters, we denote the NN function parameterized by $\theta{}\in\mathcal{R}^{N}$ as $f_{\theta{}}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{X},\mathcal{Y}$ are input and output spaces respectively. A parameter space symmetry of the NN is defined as follows (EmpiricalImpactSymmetry, ).
@@ 3.2 Exchangeable Parameters and Values in Neural Networks
In this section, we view the trained model as a random model with respect to random initializations, and thus each parameter is a random variable. We prove that certain groups of parameters $\zeta{}_{i}$’s are exchangeable, and so are values $\xi{}_{i}$’s computed with $\zeta{}_{i}$’s respectively. Below is the key insight of our proof.
Key Insight. In the initial model, $\zeta{}_{i}$’s have exchangeable distributions, as most popular NN initialization schemes draw parameters from $\mathit{iid}$ distributions. Some pose additional constraints on orthogonality or unit variances (orthogonalinit, ; unitvarianceinit, ) that introduce dependence but do not break exchangeability. Therefore, if each training step, as a transformation on the random variables, preserves exchangeability of $\zeta{}_{i}$’s, $\zeta{}_{i}$’s have exchangeable distributions in the trained model.
@@
Given an NN architecture with function $f_{\theta{}}$, assume that $\zeta{}_{i}$’s have exchangeable distributions in the initial model. If for any permutation matrix $P$, $\omega_{P}$ is a parameter space symmetry, then $\zeta{}_{i}$’s have exchangeable distributions in the trained model.
@@
Here we assume using gradient descent for supervised learning for simplicity, but the the proof easily generalizes to other optimization algorithms. As we are only interested in $\zeta{}_{i}$’s, we formalize a training step as a transformation on only the $\zeta{}_{i}$’s. [^1] In fact, one can prove that permuting $\zeta{}_{i}$’s does not affect the update to other parameters in $\theta{}^{\prime}$. Denote the NN loss function as $L_{\zeta{}}:\mathcal{X}\times\mathcal{Y}\times\mathcal{R}^{N-mn}\rightarrow\mathcal{R}$, parameterized by $\zeta{}$. $L_{\zeta{}}(x,y,\theta{}^{\prime})=\psi(f_{\theta{}}(x),y)$ for some metric function $\psi$ such as cross entropy. A training step $G_{S}:(\mathcal{R}^{m})^{n}\rightarrow(\mathcal{R}^{m})^{n}$ takes a training batch of $B$ samples $S\in(\mathcal{X}\times\mathcal{Y})^{B}$, and does a gradient descent step $G_{S}(\zeta{})=\zeta{}-\gamma\sum_{(x,y)\in S}\nabla_{\zeta{}}L_{\zeta{}}(x,y,\theta{}^{\prime})$, where $\gamma$ is the learning rate, and $\nabla_{\zeta{}}L_{\zeta{}}(x,y,\theta{}^{\prime})$ is the gradient of $L_{\zeta{}}$ with respective to $\zeta{}$, evaluated at $(x,y,\theta{}^{\prime})$.
@@
Note that $\nabla_{P\zeta{}}L_{P\zeta{}}$ is not the same as $\nabla_{\zeta{}}L_{P\zeta{}}$, as the derivatives in the gradient vector of the former case shall match the parameter order in $P\zeta{}$ rather than $\zeta{}$. Since $\omega_{P}$ is a parameter space symmetry, by definition 2, $\forall\zeta{}\in(\mathcal{R}^{m})^{n}$, $f_{\theta{}}$ and $f_{\omega_{P}(\theta{})}$ are the same function, and thus $L_{\zeta{}}$ and $L_{P\zeta{}}$ are the same function. Therefore, it is straightforward that $\nabla_{P\zeta{}}L_{P\zeta{}}(x,y,\theta{}^{\prime})=P\nabla_{\zeta{}}L_{\zeta{}}(x,y,\theta{}^{\prime})$ for all $x,y,\zeta{},\theta{}^{\prime}$. This directly leads to $PG_{S}(\zeta{})=G_{S}(P\zeta{})$. ∎
@@
Under the conditions of Theorem 2, for any NN input $x\in\mathcal{X}$, define $\xi{}_{i}=g_{\theta{}^{\prime},\zeta{}_{i}}(x)$, where $g$ is the function of a sub-network in the trained model, parameterized by one $\zeta{}_{i}$ and possibly $\theta{}^{\prime}$. We have that $\xi{}_{i}$’s have exchangeable distributions in the trained model.
@@
[…]
Note that given a group of exchangeable parameters $\zeta{}_{i}$’s, there might be multiple groups of exchangeable values $\xi{}_{i}$’s in NN computation, each with a different sub-network $g$.